OpenStudy (anonymous):

I'm having a brainfart lol. Someone please help. 4 Cylinders and 5 prisms weigh 32 oz. 1 Cylinder and 8 Prisms weigh 35 oz. How much does 1 Cylinder weigh? 1 Prism?

jimthompson5910 (jim_thompson5910):

4 Cylinders and 5 prisms weigh 32 oz. means 4c + 5p = 32 where c = weight of 1 cylinder, p = weight of 1 prism

jimthompson5910 (jim_thompson5910):

1 Cylinder and 8 Prisms weigh 35 oz. means 1c + 8p = 35

OpenStudy (anonymous):

Then just solve for each?

jimthompson5910 (jim_thompson5910):

exactly, it might help to solve for c in the second equation 1c + 8p = 35 c = -8p + 35 then plug this into the first equation 4c + 5p = 32 4(-8p + 35) + 5p = 32 to solve for p

jimthompson5910 (jim_thompson5910):

once you have p, use it to find c

OpenStudy (anonymous):